OpenStudy (anonymous):

the cost of 3 lunches and 4 dinners is $60 the cost of 5 lunches and 2 dinners is $44 Find the cost of one dinner.

OpenStudy (anonymous):

3L + 4D = $60 5L + 2D = $44 Then do simultaneous equations.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

$44-5L = 2D -> $88-10L = 4D $60-3L = 4D $88-10L = $60-3L 7L = $28 L = $4 D = ($44-5L)/2 = ($44-5($4))/2 = $24/2 = $12 Lunches are $4 and Dinners are $12